# Rigidity and nonexistence of CMC hypersurfaces in 5-manifolds

### v2

@@ Abstract.
We prove that the nonnegative $3$-intermediate Ricci curvature and uniformly positive $k$-triRic curvature implies rigidity of complete noncompact two-sided weakly stable minimal hypersurfaces in a Riemannian manifold $(X^{5},g)$ with bounded geometry. The nonnegativity of $3$-intermediate Ricci curvature can be replaced by nonnegative Ricci and biRic curvature. In particular, there is no complete noncompact stable minimal hypersurface in a closed $5$-dimensional manifold with positive sectional curvature. It extends result of Chodosh-Li-Stryker [to appear in J. Eur. Math. Soc (2024)] to $5$-dimensions. We also prove some nonexistence results of complete noncompact constant mean curvature hypersurfaces with finite index in hyperbolic space $\mathbb{H}^{5}$.
@@ 1. Introduction
An immersed complete two-sided stable minimal hypersurface $M^{n}$ in $(X^{n+1},g)$ has zero mean curvature and satisfies the inequality
(1.1) $\int_{M}\left(|A_{M}|^{2}+\operatorname{Ric}_{X}(\eta,\eta)\right)\psi^{2}\leq\int_{M}|\nabla\psi|^{2}$
for any $\psi\in C^{\infty}_{c}(M)$, where $A_{M}$ is the second fundamental form of the immersion, $\eta$ is a unit normal vector on $M$ and $\operatorname{Ric}_{X}$ is the ambient Ricci curvature. In a similar manner of stable geodesic in $2$-dimensions, stable minimal hypersurfaces is a powerful tool to study the ambient geometry of $(X^{n+1},g)$ in higher dimensions. Assuming $M$ is compact, basic arguments show that
@@
    When $\operatorname{Ric}_{X}>0$, there are no closed two-sided stable minimal hypersurfaces. If $\operatorname{Ric}_{X}\geq 0$, then $M$ is totally geodesic and $\operatorname{Ric}_{X}(\eta,\eta)\equiv 0$ along $M$; see [Simons-minimal-varieties].
@@
    If the ambient scalar curvature is strictly positive, then $M$ also admits a metric of positve scalar curvature [Schoen-Yau-incompressible-minimal-surfaces, Schoen-Yau-PSC]. In three dimension, each component of $M$ must be a $2$-sphere or projective plane.
@@
    When $\operatorname{Ric}_{X}\geq 0$, $(M,g)$ is totally geodesic, intrinsically flat and has $\operatorname{Ric}_{X}(\eta,\eta)\equiv 0$ along $M$ [Schoen-Yau-positive-Ricci-curvature]. Moreover, either $X^{3}$ is diffeomorphic to $\mathbb{R}^{3}$ or its universal cover is isometric to the Riemann product $M^{2}\times\mathbb{R}$. [Liu-nonnegative-Ricci-curvature].
@@
Recently, Chodosh, Li and Stryler developed such theory in $4$-dimensions by combining nonnegativity of $2$-intermediate Ricci curvature with strict positivity of scalar curvature.
@@
where $\operatorname{Ric}_{2}^{X}$ is the $2$-intermediate Ricci curvature (Definition 1.7). Then any complete two-sided stable minimal immersion $M^{3}\rightarrow(X^{4},g)$ is totally geodesic and has $\operatorname{Ric}_{X}(\eta,\eta)\equiv 0$ along $M$.
It is natural to ask what happens in $5$-dimensions. For example, a basic question is to determine which ambient curvature conditions imply compactness or non-existence of complete stable minimal hypersurfaces. As discussed above in $3$-manifolds, $R_{X}\geq 1$ implies compactness and $\operatorname{Ric}_{X}>0$ implies non-existence. Two examples constructed in [chodosh-li-stryker][Example 1.1,1.2] can be easily adapted to show that neither of these results can hold in $5$-dimensions.
Note that for most of aforementioned results, it is essential that no assumption is made on properness or volume growth of $M$. The result in [chodosh-li-stryker] cannot extend to $5$-dimensions directly because $\operatorname{Ric}^{X}_{2}\geq 0$ and $R_{X}>\delta$ are tailor to obtain one-endness and volume growth of hypersurfaces in $4$-dimensions. Hence, exploring alternative curvature conditions to maintain one-endness and volume growth in higher dimensions is one of the main topics in this paper.
@@ 1.1. Main results
The purpose of this article is twofold. First, we extend results in [chodosh-li-stryker] to $5$-dimensions under different curvature assumptions and weak stability of minimal hypersurface. Second, we prove that there are no complete noncompact constant mean curvature (CMC) hypersurfaces with (averaged) mean curvature $H_{M}>\frac{\sqrt{65}}{8}$, finite index and universal subexponential end growth in hyperbolic space $\mathbb{H}^{5}$, as a special case of a more general nonexistence result.
@@
Let $(X^{n+1},g)$, $n\geq 2$, be a Riemannian manifold. An immersed CMC hypersurface (two-sided if it is minimal) $M^{n}$ in $X$ is weakly stable if
$\int_{M}\left(|A_{M}|^{2}+\operatorname{Ric}_{X}(\eta,\eta)\right)\psi^{2}\leq\int_{M}|\nabla\psi|^{2}$
for $\psi\in C_{c}^{\infty}(M)$ satisfying $\int_{M}\psi=0.$ We call it stable if we do not assume $\int_{M}\psi=0$. Moreover, $M$ is said to have finite index if the operator $\Delta_{M}+|A_{M}|^{2}+\operatorname{Ric}_{X}(\eta,\eta)$ has finite number of negative Dirichlet eigenvalues in any compact subset of $M$.
@@
Let $(X^{n+1},g)$, $n\geq 2$, be a Riemannian manifold. Given a local orthonormal basis $\{e_{i}\}_{i=1}^{n+1}$ on $T_{p}X$ and $k>0$, we define $k$-weighted triRicci curvature as
(1.2) $k\operatorname{-triRic}_{X}(e_{i};e_{j},e_{l})=k\operatorname{Ric}_{X}(e_{i},e_{i})+\operatorname{biRic}_{e_{i}^{\perp}}(e_{j},e_{l}),$
where $\operatorname{biRic}^{e_{i}^{\perp}}(e_{j},e_{l})$ is given by
$\operatorname{biRic}_{e_{i}^{\perp}}(e_{j},e_{l}):=\sum_{m\neq i}\operatorname{Rm}^{X}_{mjjm}+\sum_{m\neq i}\operatorname{Rm}^{X}_{mllm}-\operatorname{Rm}^{X}_{jllj},\quad j\neq l$
i.e., we take the partial contraction on the subspace $e_{i}^{\perp}:=\operatorname{span}\{e_{l}\}_{l\neq i}$ normal to $e_{i}$.
@@
When $n=2$ or $3$, $\operatorname{1-triRic}$ is just the half of the scalar curvature of $X$. In particular, $1$-triRic curvature is the $3$-intermediate curvature $\mathcal{C}_{3}$ defined by Brendle-Hirsch-Johne[brendlegeroch'sconjecture].
@@
$\begin{split}&\sum_{i=1}^{n+1}\sum_{j\neq i,l\neq i}k\operatorname{-triRic}_{X}(e_{i};e_{j},e_{l})\\ &=\sum_{i=1}^{n+1}\left(kn(n-1)\operatorname{Ric}_{X}(e_{i},e_{i})+(2n-3)\sum_{j\neq i,l\neq i}\operatorname{Rm}^{X}_{jllj}\right)\\ &=kn(n-1)\sum_{i=1}^{n+1}\operatorname{Ric}_{X}(e_{i},e_{i})+(2n-3)(n-1)\sum_{j,l=1}^{n+1}\operatorname{Rm}^{X}_{jllj}\\ &=((k+2)n-3)(n-1)R_{X}.\end{split}$
In particular, positive $k$-weighted triRicci curvature implies that $X$ has positive scalar curvature. On the converse, $\Sigma^{2}\times S^{3}(5/12)$ has positive scalar curvature but not always positive $1$-weighted triRicci curvature where $\Sigma^{2}$ is a closed hyperbolic surface and $S^{3}(5/12)$ is a sphere with constant curvature $5/12$.
@@
Let $(X^{n+1},g)$, $n\geq 2$, be a Riemannian manifold. The $(n-1)$-intermediate Ricci curvature of $X$ in the direction $\mu$, is defined as
$\operatorname{Ric}^{X}_{n-1}(e_{i};\mu)=\sum_{j\in\{1,\cdots,n\}\setminus i}\operatorname{Rm}_{X}(e_{j},\mu,\mu,e_{j})$
where $\{e_{1},\cdots,e_{n},\mu\}$ is any orthonormal basis of $T_{p}X$.
Throughout this article, for any tensor $T$ acting on tangent vectors we say $T\geq C$ for a constant $C$ if $T(e_{i},\cdots)\geq C$ for any unit vectors $e_{i},\cdots\in T_{p}X$ and any $p\in X.$ Note that $\operatorname{Ric}_{n-1}^{X}\geq 0$ implies that $1\operatorname{-triRic}_{X}\geq 0$.
We now turn to our first main result.
@@
$\operatorname{Ric}^{X}_{3}\geq 0\ \text{and}\ k\operatorname{-triRic}_{X}\geq\delta>0\ \text{for}\ k\in[3/2,2].$
@@
Then any complete two-sided weakly stable minimal immersion $M^{4}\rightarrow X^{5}$ is totally geodesic and has $\operatorname{Ric}_{X}(\eta,\eta)\equiv 0$ along $M.$
It is possible to weaken non-negativity of $\operatorname{Ric}_{3}^{X}$ to a mixed curvature condition. Note that $\operatorname{Ric}_{3}^{X}\geq 0$ implies $\operatorname{Ric}_{X}\geq 0$ and $\operatorname{biRic}_{X}\geq 0$, where $\operatorname{biRic}_{X}$ is the bi-Ricci curvature (see Definition 2.2 in detail). In the Appendix, we provide a (one parameter family) closed $5$-dimensional Riemannian manifold $(X^{5},g)$, satisfying
$\operatorname{Ric}_{X}>0,\ \,\operatorname{biRic}_{X}>0,\ \,\text{and}\ \ 2\operatorname{-triRic}_{X}\geq\delta>0,$
while $\operatorname{Ric}_{3}<0$ somewhere. As a matter of fact, the result in Theorem 1.8 holds with $\operatorname{Ric}_{3}^{X}\geq 0$ replaced by $\operatorname{Ric}_{X}\geq 0$ plus $\operatorname{biRic}_{X}\geq 0$.
Theorem 1.9.
$\operatorname{Ric}_{X}\geq 0,\ \,\operatorname{biRic}_{X}\geq 0,\ \,\text{and}\ \ k\operatorname{-triRic}_{X}\geq\delta>0\ \text{for}\ k\in[3/2,2].$
Then any complete two-sided weakly stable minimal immersion $M^{4}\rightarrow X^{5}$ is totally geodesic and has $\operatorname{Ric}_{X}(\eta,\eta)\equiv 0$ along $M.$
@@
Let us make some comments here. The first one is that if $(X^{5},g)$ is a manifold with bounded geometry satisfying $\operatorname{Ric}_{X}>0$, $\operatorname{biRic}_{X}\geq 0$ and $k\operatorname{-triRic}_{X}\geq\delta>0\ \text{for}\ k\in[3/2,2]$, then there is no complete two-sided weakly stable minimal immersion. This follows directly from Theorem 1.9. The second one is that there is no complete two-sided weakly stable minimal immersion in manifold $X$ with $\operatorname{Ric}_{X}>0$ and $\operatorname{biRic}_{X}\geq\delta>0$. Indeed, uniformly positive $\operatorname{biRic}_{X}$ implies that a complete two-sided weakly stable minimal immersion must be compact [shenyingyerugang]. Then the standard argument leads to contradiction. The third one is that Theorem 1.9 also implies that there is no two-sided weakly stable minimal immersion in $(\mathbb{S}^{5},g_{can})$ (previously proved by [catino]). Lastly, the proof of above theorems also work for weakly stable CMC immersion.
The second result is about the nonexistence of complete noncompact CMC hypersurfaces with finite index. We first recall universal subexponential end growth condition for a noncompact manifold $M$ (possibly with noncompact boundary).
Definition 1.10 ([Hong24]).
For a fixed $R>0$ and $p\in M$, define $N_{M,k}$ to be the number of components of $M\cap(B_{(k+1)L+R}\setminus B_{kL+R})$:
$N_{M,k}:=\#\{M\cap(B_{(k+1)L+R}\setminus B_{kL+R})\}.$
We say that $M$ has subexponential end growth if
$\limsup_{k\rightarrow\infty}\frac{N_{M,k}}{e^{Ck}}=0$
for any constant $C>0.$
Note that if $M$ has boundary, $M\cap(B_{(k+1)L+R}\setminus B_{kL+R})$ contains two types of components, depending on whether they have boundary portion belonging to $\partial M.$
Definition 1.11 ([Hong24]).
A complete noncompact manifold $M$ has universal subexponential end growth if there is one connected component of $M\setminus B_{R_{0}}$ whose universal covering satisfies subexponential end growth for some fixed $R_{0}>0$.
Theorem 1.12.
There is no complete noncompact constant mean curvature hypersurface $M$ in hyperbolic space $\mathbb{H}^{5}$ satisfying universal subexponential end growth condition,
$H_{M}>\frac{\sqrt{65}}{8}\ \ \text{and}\ \ \operatorname{Ind}<\infty.$
The question of whether a complete constant mean curvature hypersurface in $\mathbb{H}^{n+1}$, where $n\geq 4$, with finite index and $H>1$ must be compact remains open in general. Previously, Da Silveira confirmed this for the case $n+1=3$ in [daSilveira-Stability-of-complete], while the positive answer for $n+1=4$ was provided by the first author in [Hong24] under the universal subexponential end growth condition. Partial progress has been made under various stronger or additional assumptions. For instance, in dimension $n+1=5$ [chengxufiniteindex, dengqintao], an assumption was imposed on mean curvature: $H>\frac{\sqrt{175}}{\sqrt{148}}$. Thus, our theorem presents a contribution when $H\in(\frac{\sqrt{65}}{8},\frac{\sqrt{175}}{\sqrt{148}}]$. We remark that the universal subexponential end growth condition in above theorem can be replaced by finite first Betti number and finite ends. Let us mention that when zero volume entropy is additionally assumed, a complete positive solution was achieved in all dimensions [zeroentropy].
Indeed, we can prove a nonexistence result under more general settings; see Theorem 5.2 in detail. Consequently, Theorem 1.12 follows from Theorem 5.2 directly.

### v3

# Rigidity of CMC hypersurfaces in 5-and 6-manifolds

@@ Abstract.
We prove that nonnegative $3$-intermediate Ricci curvature combined with uniformly positive $k$-triRic curvature implies rigidity of complete noncompact two-sided stable minimal hypersurfaces in a Riemannian manifold $(X^{5},g)$ with bounded geometry. The stonger assumption of nonnegative $3$-intermediate Ricci curvature can be replaced by the nonnegativity of Ricci and biRic curvature. In particular, there is no complete noncompact stable minimal hypersurface in a closed $5$-dimensional manifold with positive sectional curvature. This extends result of Chodosh-Li-Stryker [J. Eur. Math. Soc (2025)] to $5$-dimension. We also establish rigidity results on CMC hypersurfaces with nonzero mean curvature in $5$- and $6$-manifolds.
@@
CMC hypersurfaces, Rigidity
2020 Mathematics Subject Classification:
Primary 53A10
@@ 1. Introduction
An immersed complete two-sided stable minimal hypersurface $M^{n}$ in orientable manifold $(X^{n+1},g)$ has zero mean curvature and satisfies the inequality
(1.1) $\int_{M}\left(|A_{M}|^{2}+\operatorname{Ric}^{X}(\eta,\eta)\right)\psi^{2}\leq\int_{M}|\nabla\psi|^{2}$
for any $\psi\in C^{\infty}_{c}(M)$, where $A_{M}$ is the second fundamental form of the immersion, $\eta$ is a unit normal vector on $M$ and $\operatorname{Ric}^{X}$ is the ambient Ricci curvature tensor. In a similar manner of stable geodesic in $2$-dimensions, stable minimal hypersurface is a powerful tool to study the ambient geometry of $(X^{n+1},g)$ in higher dimensions. Assuming $M$ is compact and orientable, standard arguments show that
@@
    When $\operatorname{Ric}^{X}>0$, there are no closed two-sided stable minimal hypersurfaces. If $\operatorname{Ric}^{X}\geq 0$, then $M$ is totally geodesic and $\operatorname{Ric}^{X}(\eta,\eta)\equiv 0$ along $M$; see [Simons-minimal-varieties].
@@
    If the ambient scalar curvature is strictly positive, then $M$ also admits a metric of positive scalar curvature [Schoen-Yau-incompressible-minimal-surfaces, Schoen-Yau-PSC]. In three dimension, each component of $M$ must be a $2$-sphere.
@@
    When $\operatorname{Ric}^{X}\geq 0$, $(M,g)$ is totally geodesic, intrinsically flat and has $\operatorname{Ric}^{X}(\eta,\eta)\equiv 0$ along $M$ [Schoen-Yau-positive-Ricci-curvature]. Moreover, either $X^{3}$ is diffeomorphic to $\mathbb{R}^{3}$ or its universal cover is isometric to the Riemann product $M^{2}\times\mathbb{R}$ [Liu-nonnegative-Ricci-curvature].
@@
Recently, Chodosh, Li and Stryker developed such theory in $4$-dimensions by combining nonnegativity of $2$-intermediate Ricci curvature with strict positivity of scalar curvature.
@@
where $\operatorname{Ric}_{2}^{X}$ is the $2$-intermediate Ricci curvature (Definition 1.7). Then any complete two-sided stable minimal immersion $M^{3}\rightarrow(X^{4},g)$ is totally geodesic and has $\operatorname{Ric}^{X}(\eta,\eta)\equiv 0$ along $M$.
It is natural to ask what happens in higher dimensions. A fundamental question is to determine which ambient curvature conditions ensure compactness or rule out the existence of complete stable minimal hypersurfaces. As discussed above in $3$-manifolds, $R_{X}\geq 1$ implies compactness and $\operatorname{Ric}^{X}>0$ implies non-existence. However, two examples constructed in [chodosh-li-stryker][Example 1.1,1.2] can be easily adapted to show that neither of these results can hold in higher dimensions.
Note that for most of aforementioned results, it is essential that no assumption is made on properness or volume growth of $M$. The result in [chodosh-li-stryker] cannot extend to $5$-dimensions directly because $\operatorname{Ric}^{X}_{2}\geq 0$ and $R_{X}>\delta$ are tailor to obtain one-endness and volume growth of hypersurfaces in $4$-dimensions. Hence, exploring alternative curvature conditions to maintain one-endness and volume growth in higher dimensions is the main topic in this paper.
@@ 1.1. Main results
The purpose of this article is to extend results in [chodosh-li-stryker] to $5$ and $6$ dimensions under different curvature assumptions and stability of CMC hypersurfaces.
@@
Let $(X^{n+1},g)$, $n\geq 2$, be a Riemannian manifold. An immersed CMC hypersurface (two-sided if it is minimal) $M^{n}$ in $X$ is stable if
$\int_{M}\left(|A_{M}|^{2}+\operatorname{Ric}^{X}(\eta,\eta)\right)\psi^{2}\leq\int_{M}|\nabla\psi|^{2}$
for $\psi\in C_{c}^{\infty}(M)$.
@@
Let $(X^{n+1},g)$, $n\geq 2$, be a Riemannian manifold. Given a local orthonormal basis $\{e_{i}\}_{i=1}^{n+1}$ of $T_{p}X$ and $k,\alpha>0$, we define $(k,\alpha)$-triRicci curvature in the direction $(e_{i};e_{j},e_{l})$ as
(1.2) $(k,\alpha)\operatorname{-triRic}_{X}(e_{i};e_{j},e_{l}):=k\operatorname{Ric}^{X}(e_{i},e_{i})+\alpha\text{-}\operatorname{biRic}_{e_{i}^{\perp}}(e_{j},e_{l}),$
where $\alpha\text{-}\operatorname{biRic}_{e_{i}^{\perp}}(e_{j},e_{l})$ is given by
$\alpha\text{-}\operatorname{biRic}_{e_{i}^{\perp}}(e_{j},e_{l}):=\sum_{m\neq i}\operatorname{R}^{X}_{mjmj}+\alpha\sum_{m\neq i}\operatorname{R}^{X}_{mlml}-\alpha\operatorname{R}^{X}_{jljl},\quad j\neq l$
i.e., we take the partial contraction on the subspace $e_{i}^{\perp}:=\operatorname{span}\{e_{l}\}_{l\neq i}$ normal to $e_{i}$. Furthermore, we define
$(k,\alpha)\operatorname{-triRic}_{X}(p)=\inf\{(k,\alpha)\operatorname{-triRic}_{X}(e_{i};e_{j},e_{l})\},$
where the infimum is taken over all local orthonormal basis $\{e_{i}\}_{i=1}^{n+1}$ at $p$, and over all admissible index triples $(e_{i};e_{j},e_{l})$, $i\neq j\neq l$. If $\alpha=1$, we simply denote it by $k\operatorname{-triRic}_{X}$.
@@
When $n=2$ or $3$, $1\operatorname{-triRic}_{X}$ is just the half of the scalar curvature of $X$. In higher dimensions, $1\operatorname{-triRic}_{X}$ is the $3$-intermediate curvature $\mathcal{C}_{3}$ defined by Brendle-Hirsch-Johne [brendlegeroch'sconjecture].
@@
$\begin{split}&\sum_{i=1}^{n+1}\sum_{j\neq i,l\neq i}k\operatorname{-triRic}_{X}(e_{i};e_{j},e_{l})\\ &=\sum_{i=1}^{n+1}\left(kn(n-1)\operatorname{Ric}^{X}(e_{i},e_{i})+(2n-3)\sum_{j\neq i,l\neq i}\operatorname{R}^{X}_{jljl}\right)\\ &=kn(n-1)\sum_{i=1}^{n+1}\operatorname{Ric}^{X}(e_{i},e_{i})+(2n-3)(n-1)\sum_{j,l=1}^{n+1}\operatorname{R}^{X}_{jljl}\\ &=((k+2)n-3)(n-1)R^{X}.\end{split}$
In particular, positive $k$-triRicci curvature implies that $X$ has positive scalar curvature. Conversely, $\Sigma^{2}\times S^{3}(5/12)$ has positive scalar curvature but does not have positive $1$- triRicci curvature where $\Sigma^{2}$ is a closed hyperbolic surface and $S^{3}(5/12)$ is a sphere with constant curvature $5/12$.
@@
Let $(X^{n+1},g)$, $n\geq 2$, be a Riemannian manifold. The $(n-1)$-intermediate Ricci curvature of $X$ at $p\in X$ in the direction $e_{l}$, is defined as
$\operatorname{Ric}^{X}_{n-1}(e_{i};e_{l})=\sum_{j\in\{1,\cdots,n\}\setminus i}\operatorname{R}_{X}(e_{j},e_{l},e_{j},e_{l})$
where $\{e_{1},\cdots,e_{n}\}$ is any orthonormal basis of $T_{p}X$. Similarly, we define
$\operatorname{Ric}^{X}_{n-1}(p)=\inf\{\operatorname{Ric}^{X}_{n-1}(e_{i};e_{l}):\ i\neq l\}.$
We now state our results on minimal hypersurfaces in 5-manifolds; more general results can be found in Section 5.
Theorem 1.8.
If $(X^{5},g)$ is a complete $5$-dimensional Riemannian manifold with weakly bounded geometry, satisfying
$\operatorname{Ric}^{X}_{3}\geq 0\ \text{and}\ k\operatorname{-triRic}_{X}\geq\delta>0\ \text{for}\ k\in[1,2].$
Then any complete two-sided stable minimal immersion $M^{4}\hookrightarrow X^{5}$ is parabolic with at most two ends. In particular, it is totally geodesic and has $\operatorname{Ric}^{X}(\eta,\eta)\equiv 0$ along $M.$
It is possible to weaken non-negativity of $\operatorname{Ric}_{3}^{X}$ to a mixed curvature condition. Note that $\operatorname{Ric}_{3}^{X}\geq 0$ implies $\operatorname{Ric}^{X}\geq 0$ and $\operatorname{biRic}_{X}\geq 0$, where $\operatorname{biRic}_{X}$ is the bi-Ricci curvature (see Definition 2.2 in detail). In the Appendix, we provide a (one parameter family) closed $5$-dimensional Riemannian manifold $(X^{5},g)$, satisfying
$\operatorname{Ric}^{X}>0,\ \,\operatorname{biRic}_{X}>0,\ \,\text{and}\ \ 1\operatorname{-triRic}_{X}\geq\delta>0,$
while $\operatorname{Ric}^{X}_{3}<0$ somewhere. As a matter of fact, the result in Theorem 1.8 holds with $\operatorname{Ric}_{3}^{X}\geq 0$ replaced by $\operatorname{Ric}^{X}\geq 0$ plus $\operatorname{biRic}_{X}\geq 0$.
@@
$\operatorname{Ric}^{X}\geq 0,\ \,\operatorname{biRic}_{X}\geq 0,\ \,\text{and}\ \ k\operatorname{-triRic}_{X}\geq\delta>0\ \text{for}\ k\in[1,2].$
@@
Then any complete two-sided stable minimal immersion $M^{4}\rightarrow X^{5}$ is parabolic with at most two ends. In particular, it is totally geodesic and has $\operatorname{Ric}^{X}(\eta,\eta)\equiv 0$ along $M.$
@@
Let us make some comments here. The first one is that if $(X^{5},g)$ is a manifold with bounded geometry satisfying $\operatorname{Ric}^{X}>0$, $\operatorname{biRic}_{X}\geq 0$ and $k\operatorname{-triRic}_{X}\geq\delta>0\ \text{for}\ k\in[1,2]$, then there is no complete two-sided stable minimal immersion. This follows directly from Theorem 1.9. As a consequence, there is no two-sided stable minimal immersion in $(\mathbb{S}^{5},g_{can})$ (previously proved in [catino]). The second one is that there is no complete two-sided stable minimal immersion in manifold $X$ with $\operatorname{Ric}^{X}>0$ and $\operatorname{biRic}_{X}\geq\delta>0$. Indeed, uniformly positive $\operatorname{biRic}_{X}$ implies that a complete two-sided stable minimal immersion must be compact [shenyingyerugang]. Then the standard argument leads to contradiction. Finally, we would like to mention that the conclusion of at most two ends in above two theorems can be compared to the result proved by Fischer-Colbrie and Schoen [Fischer-Colbrie-Schoen-The-structure-of-complete-stable] that a complete two-sided stable minimal surface in a manifold with nonnegative scalar curvature has at most two ends (plane or cylinder). Furthermore, we prove some related results in dimension $n+1=6$ (see Section 5.2).